REU: SCHEDULE FOR THE FIRST TWO WEEKS

There will be an opening meeting Monday, June 23, at 9:00.

Pizza and coke will be available around noon that day.

YSP/SESAME prepraration and training sessions will take place Monday 10:00 – 12:00 and Tuesday through Friday 9:00 – 12:00.

Date  YSP/SESAME  Babai  Farb  Fefferman Herrmann
Monday, June 23  10:00 – 12:00 3:30 – 6:00  1:00 – 3:00 7:00 – 8:30
Tuesday, June 24  9:00 – 12:00 3:30 – 6:00 1:00 – 3:00  7:00 – 8:30
Wednesday, June 25 9:00 – 12:00 3:30 – 6:00  1:00 – 3:00 7:00 – 8:30
Thursday, June 26  9:00 – 12:00 3:30 – 6:00 1:00 – 3:00  7:00 – 8:30
Friday, June 27  9:00 –12:00 3:30 – 6:00  1:00 – 3:00
Date  YSP/SESAME  Babai  Farb  Kirr  Herrmann
Monday, June 30  9:30 – 2:30 2:45 – 4:00  7:00 – 8:30 4:15 – 5:30
Tuesday, July 1  9:30 – 2:30  2:45 – 4:00 7:00 – 8:30 4:15 – 5:30
Wednesday, July 2  9:30 – 2:30 2:45 – 4:00  7:00 – 8:30 4:15 – 5:30
Thursday, July 3  9:30 – 2:30  2:45 – 4:00

Graduate student sessions will be organized around this schedule.

Here is the schedule in summary form:

First week:

YSP/SESAME 9:00 – 12:00 M – F Babai 3:30 – 6:00 M – F Farb 1:00 – 3:00 TTh Fefferman 1:00 – 3:00 MWF Herrmann 7:00 – 8:30 M – Th

Second week:

YSP 9:30 – 2:30 M – Th Babai 2:45 – 4:00 MW Farb 2:45 – 4:00 TTh Herrmann 4:15 – 5:30 M – W Kirr 7:00 – 8:30 M – W